Technology. Driving down a mountain, Tom finds that he has descended 1800 ft
in elevation by the time he is 3.25 mi horizontally away from the top of the
mountain. How do I find the slope of his descent to the nearest hundredth?

Answer by checkley75(3666)

using the graphing technique we TWO POINTS (0,1800) & (3.25*5280,0)
OR (0,1800) & (17160,0)
NOW WE FIND THE SLOPE BETWEEN THESE TWO POINTS THUS:
(0-1800)/(17160-0)=-1800/17600=-.1022727 FOR THE SLOPE OR (.10).